2. Assume that the price index (the measure of the economy's price levels in the current time is 1.35.


A. If you are to compare the prices of the current time and the previous time (the previous time being the reference period), how much is the change in the economy's price level?

B. Determine the value of peso at the current time period. Interpret the result.

C. Determine exactly how much (in percentage) the did the value of peso decreased.

Interpret the result.

D. If, instead, the price level falls to 0.80:

D.1 how much is the change in the economy's price level?

D.2 what is the value of the peso at the current time

period? Interpret the result.

D.2 determine exactly how much (in percentage) the did the value of peso change.


1
Expert's answer
2021-03-04T10:35:48-0500
"solution"

"given;\\
price\ in\ current\ time=1.35\\
price\ level\ in\ the\ base\ period\ is\ always=1.00\\"


A

"\% change \ in\ price\ level\ \\
= \frac {1.35-1.00}{1.00}=0.35=35\%"

B

"value\ of\ peso\ at\ the\ current\ time\ period\\
=\frac{1}{1.35}\\"


interpretation: if one peso could buy one unit of a good"\ K," then at the current period one peso can buy "\frac {1}{1.35}" units of "\ K," other things being constant .


C

"\%change\ in\ value\ of \ peso\\
=[\frac{0.74074-1}{1}]=0.25926=25.92\%"

decrease in value of peso is equal to "25.92\%" , that is if a person with some income can now only buy "74.08\%" of the basket of good that he could buy in the base period with the same income.

D

"price\ level\ in \ current \ time=0.80\\
[i ] \% change \ in \ price\ level \\
\frac{0.80-1.00}{1.00} =-0.2=-20\%\\
therefore\ there\ is\ deflation\ of \ 20\%"


"[ii]value\ of current\ time\ priod\ of\ pesos\\
=[\frac {1}{0.80}]=1.25\\"


that is if one peso could buy one unit of some good "\ X" in base period ,then in the current period one peso can buy "1.25" units of the same goods all being constant .


"[iii]\%change \ in \ value\ of \ peso \\
=[\frac {1.25-1}{1}]=0.25=25\%\\
therefore \ the \ value\ of \ peso\ increased \ by\ 25\%"